OpenStudy (anonymous):

Calc I: lim(fg) = lim f lim g but (fg) ' not = f ' g ' Shouldn't the above be equivalent since f ' = lim f ?

OpenStudy (turingtest):

lim f has no meaning unless you specify what the limit is approaching

OpenStudy (turingtest):

but either way lim f is not the same as f'

OpenStudy (turingtest):

so what are yo asking for, the proof of the product rule?

OpenStudy (turingtest):

you must remember that\[f'(x)=\lim_{h\to0}{f(x+h)-f(x)\over h}\]that is not just any old limit, but the derivative...

OpenStudy (anonymous):

My teaches told us that 'derivative' is just another name for limit. Sum and constant multiple rules seem to match, but then product and power rules don't.... Still a bit confused, but tnx for help

OpenStudy (turingtest):

your teacher sucks then, sorry

OpenStudy (turingtest):

I will prove the product rule for you if you have the patience, but @Algebraic! may be doing it as we speak

OpenStudy (turingtest):

...or not

OpenStudy (anonymous):

I think we're assuming we're talking about the limit as the definition of the derivative.

OpenStudy (turingtest):

right, but surely you agree that the limit and derivative are not names for the same thing; that is terrible terminology at best

OpenStudy (turingtest):

the limit of a function as x approaches a is\[\lim_{x\to a}f(x)\]the derivative of a function at x=a is\[f'(a)=\lim_{h\to a}{f(x)-f(a)\over x-a}\]very different things.... you want the proof or not?

OpenStudy (anonymous):

if you dont mind or maybe a link will do TNX

OpenStudy (turingtest):

nah I'll do it :)

OpenStudy (anonymous):

hugs and kisses...

OpenStudy (turingtest):

The general formula for the derivative of a function by definition is\[\frac d{dx}f(x)=\lim_{h\to0}{f(x+h)-f(x)\over h}\]Now we want to know\[\frac d{dx}[f(x)g(x)]\]so we apply the same definition...\[\frac d{dx}[f(x)g(x)]=\lim_{h\to0}{f(x+h)g(x+h)-f(x)g(x)\over h}\]Now a trick...\[=\lim_{h\to0}{f(x+h)g(x+h)-f(x)g(x+h)+f(x)g(x+h)-f(x)g(x)\over h}\]\[=\lim_{h\to0}{[f(x+h)-f(x)]g(x+h)+f(x)[g(x+h)-g(x)]\over h}\]\[=\lim_{h\to0}{f(x+h)-f(x)\over h}g(x+h)+f(x)\lim_{x\to0}{g(x+h)-g(x)\over h}\]\[\frac d{dx}[f(x)g(x)]=f'(x)g(x)+f(x)g'(x)~~~~~~~~~~~~~~~~~~~~\huge\checkmark \]

OpenStudy (turingtest):

I'm a perfectionist, so I cleaned it up no typose though, so there you have it :)

OpenStudy (turingtest):

typos* lol

OpenStudy (anonymous):

TNX again

OpenStudy (turingtest):

welcome!